Vertify identity
1+sin(data)/sin(data)=cot^2/csc(data)-1

Answer by lwsshak3(11628)

Vertify identity
1+sin(data)/sin(data)=cot^2/csc(data)-1
(1+sin)/sin=cot^2/csc-1
The right-hand side looks more complicated, so let us start with that side
(right-hand side)=cot^2/csc-1=(cos^2/sin^2)/(1/sin)-1
=(cos^2/sin^2)/((1-sin)/sin)
=cos^2/sin(1-sin)
=1-sin^2/sin(1-sin)
=(1+sin)(1-sin)/sin(1-sin)
=(1+sin)/sin=(left-hand side)
verified: right-hand side = left-hand side